# Contrastive Localized Language-Image Pre-Training

### v1

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has been a celebrated method for training vision encoders to generate image/text representations facilitating various applications. Recently, CLIP has been widely adopted as the vision backbone of multimodal large language models (MLLMs) to connect image inputs for language interactions. The success of CLIP as a vision-language foundation model relies on aligning web-crawled noisy text annotations at image levels. Nevertheless, such criteria may become insufficient for downstream tasks in need of fine-grained vision representations, especially when region-level understanding is demanding for MLLMs. In this paper, we improve the localization capability of CLIP with several advances. We propose a pre-training method called Contrastive Localized Language-Image Pre-training (CLOC) by complementing CLIP with region-text contrastive loss and modules. We formulate a new concept, promptable embeddings, of which the encoder produces image embeddings easy to transform into region representations given spatial hints. To support large-scale pre-training, we design a visually-enriched and spatially-localized captioning framework to effectively generate region-text pseudo-labels at scale. By scaling up to billions of annotated images, CLOC enables high-quality regional embeddings for image region recognition and retrieval tasks, and can be a drop-in replacement of CLIP to enhance MLLMs, especially on referring and grounding tasks.
@@ 1 Introduction
Large-scale vision-language (VL) pre-training has been an important foundation for the recent tremendous growth of multimodal applications. Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021; Jia et al., 2021) has become a great success of VL representation learning that connects images and text by contrastive training on web-crawled image-text pairs. It has been proven strong transferability and generalizability on extensive downstream tasks such as zero-shot image classification and image-text retrieval. Even beyond, CLIP has become arguably the default choice of vision backbone for multimodal large language models (MLLMs) (Liu et al., 2023; McKinzie et al., 2024; Zhang et al., 2024a) due to its superior prior knowledge in aligning vision and language (Tong et al., 2024), facilitating vision inputs to be injected into language models.
@@
Figure 1: Overview of our CLOC pre-training framework. (1) A visually-enriched and spatially-localized captioning pipeline generates pseudo-labeled bounding boxes with detailed descriptions for key image regions. (2) A lightweight Prompter attached on top of the CLIP image encoder can be prompted to transform the image embedding into the region-focused feature. All parameters are trained end-to-end from scratch with our contrastive localized language-image loss on the annotated region-text datasets. After pre-training, (3a) region features can be generated via the Prompter for region-text tasks like object classification in a training-free fashion. (3b) The image encoder, along with the optional Prompter, can also strengthen MLLMs fine-tuning by enhancing their fine-grained image understanding capabilities.
@@
As VL research gets increasing attention and expedites progress, various more advanced multimodal tasks are demanding stronger vision capabilities. For instance, recent MLLMs (Rasheed et al., 2024; Ren et al., 2024; Lai et al., 2023; Chen et al., 2023; Peng et al., 2023) have been focusing on more fine-grained understanding tasks that require comprehension of the semantic at region levels such as visual question answering (VQA) with referring and grounding instructions. These MLLMs are fine-tuned on referring and grounding data with CLIP as the vision backbone, as seen in works like Kosmos-2 (Peng et al., 2023) and Ferret (You et al., 2023; Zhang et al., 2024b). Due to the need for such region-level understanding, CLIP, which aligns entire images with text captions, seems insufficient, as its regular image-text contrastive loss primarily emphasizes global semantics.
@@
To remedy such core localization capability for CLIP, we ask a challenging and fundamental question: without sacrificing CLIP’s original strong image-level knowledge, can we pre-train a stronger image encoder with enhanced localization capability that can be inherently integrated into MLLMs?
@@
To this end, we explore a data-driven approach that complements the original CLIP image-text pre-training objective with explicit region-text supervision. Though conceptually simple, several challenges exist. First, it lacks public datasets with region-text annotations at scales large enough for CLIP training, which typically requires hundreds of millions even billions of images. Existing region-text corpus like Visual Genome (Krishna et al., 2017) contains about $108$K images, and the largest noisy-labeled grounded dataset GRIT (Peng et al., 2023) features only $20$M images. Indeed, such deficiency has probably limited the literature to mainly consider semi-supervised or weakly-supervised approaches as somewhat a compromise (Naeem et al., 2023; Yao et al., 2022; 2023).
@@
  • We design a large-scale pseudo-labeling data engine to support CLOC training. We properly combine visual-enriched image captioners and open-vocabulary detectors for an effective recipe that improves previous practice of region annotations (Minderer et al., 2024; Peng et al., 2023). This approach yields a two-billion image-text dataset with fine-grained region-text annotations, which serves as the foundation for training our CLOC model.
@@
  • We are working on releasing our pre-trained checkpoints and the constructed region-text annotations along with the final version to accelerate future research within the community.
@@ 2 Related Work
Since CLIP was introduced, many follow-up works have been proposed to improve it from various aspects, for different target tasks, and with different approaches. From the aspect relevant to our work, improving the localization capability, most works specifically focus on the downstream dense vision tasks such as open-vocabulary detection (Minderer et al., 2024; Yao et al., 2022; Wu et al., 2023). Another less and arguably more challenging thread is to maintain the generalizability of CLIP on image-level tasks while improving localization. Recent works like SILC (Naeem et al., 2023) and SPARC (Bica et al., 2024) combine localization-enhancing unsupervised objectives with the CLIP loss, but do not attempt with supervision on large-scale explicit pseudo-labeled data like ours. Alpha-CLIP (Sun et al., 2024) shows that the SAM segmentation model (Kirillov et al., 2023) can provide useful conditions for CLIP.
Vision encoder pre-training for MLLMs.
Building upon the success of large language models (LLMs), a popular approach to MLLMs like LLaVA (Liu et al., 2023), typically connects a vision encoder (e.g., ViT (Dosovitskiy et al., 2021)) to digest visual inputs and maps them to the LLM decoder input space as token embeddings. Among various types of vision encoders (Oquab et al., 2023; He et al., 2022), CLIP (Radford et al., 2021; Jia et al., 2021) becomes the most popular choice, due to its superior performance on MLLM benchmarks reported by recent studies (Tong et al., 2024).
@@
Large-scale training data are the fuel of pre-training, especially for CLIP. The literature has been exploring scalable ways to generate high-quality synthetic annotations. For instance, several works demonstrate that visually-enriched image captions improve CLIP (Lai et al., 2024). MOFI (Wu et al., 2024) constructs a large alt-text set and augments CLIP with a multi-classification task. However, these works only consider image-level annotations but not explicit region-level labels. In the context of dense vision tasks like open-vocabulary detection and segmentation, pseudo-labeling in a self-training paradigm has proven an effective approach (Kirillov et al., 2023; Minderer et al., 2024). We are inspired by these efforts and build on them to enhance CLIP’s localization capabilities.
3 CLOC: Contrastive Localized Language-Image Pre-Training
3.1 Preliminary: from Image-Text to Region-Text Alignment
Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) trains a pair of image and text encoders (denoted as $f_{I}$ and $f_{T}$, respectively) by contrastively aligning the image and text embeddings. Let a mini-batch of $N$ image-text pairs $\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ be sampled from the large-scale training set during each training iteration. The contrastive loss is defined as follows:
${\mathcal{L}}_{\text{CLIP}}:=({\mathcal{L}}_{I\rightarrow T}+{\mathcal{L}}_{T\rightarrow I})/2.\>{\mathcal{L}}_{I\rightarrow T}:=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\big{(}\texttt{sim}(f_{I}({\bm{x}}_{i}),f_{T}({\bm{y}}_{i}))/\tau\big{)}}{\sum_{j=1}^{N}\exp\big{(}\texttt{sim}(f_{I}({\bm{x}}_{i}),f_{T}({\bm{y}}_{j}))/\tau\big{)}},$ (1)
where $\texttt{sim}(\cdot,\cdot)$ is the similarity measurement function and $\tau$ is a (learnable) logit temperature. The CLIP loss ${\mathcal{L}}_{\text{CLIP}}$ averages the symmetrical contrastive loss in which cross-entropy normalized along image-to-text and text-to-image axes, respectively.
Conceptually, the CLIP loss aligns images with their associated text, but it overlooks regional information and spatial semantics. We propose augmenting this with region-text alignment on top of ${\mathcal{L}}_{\text{CLIP}}$. Specifically, assume an image-text pair $({\bm{x}},{\bm{y}})$ can be decomposed into image regions ${\bm{x}}^{(1)},\dots,{\bm{x}}^{(m)}$, and there exist fine-grained captions ${\bm{y}}^{\prime(m)}$ that describe the corresponding image regions ${\bm{x}}^{(m)}$. Thus, the original input $({\bm{x}},{\bm{y}})$ becomes $\{({\bm{x}}^{(1)},{\bm{y}}^{\prime(1)}),\dots,({\bm{x}}^{(m)},{\bm{y}}^{\prime(m)})\}$ for region-text considerations, and $({\bm{x}},{\bm{y}})$ is a special case when the “region” itself is the whole image. Based on this, we identify several research questions and will answer them in the following sections:
@@ 3.2 Promptable Embeddings
We re-formulate the CLIP loss based on image-text pairs $({\bm{x}},{\bm{y}})$ into a localized language-image contrastive loss for region-text alignment based on triplets of ($(\{{\bm{l}}\},{\bm{x}},{\bm{y}})$, where ${\bm{l}}$ is a location representation such as a bounding box, and possibly there are several boxes as a set $\{{\bm{l}}\}$ per image. To make it compatible with CLIP training, we construct a promptable embedding transform module, or in short, region prompter ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}({\bm{x}}))$, that extracts the region embedding specified by ${\bm{l}}$ from the image embedding $f_{I}({\bm{x}})$. This formulation is inspired by the success of the segmentation model SAM (Kirillov et al., 2023) that predicts the segmentation masks conditioned on location prompt (e.g., a box), while CLOC predicts a region embedding conditioned on ${\bm{l}}$ instead.
@@
${\mathcal{L}}_{R\rightarrow T}:=-\frac{1}{MN}\sum_{i=1}^{N}\sum_{{\bm{l}}_{i}^{(m)}\sim\{{\bm{l}}_{i}\}}\log\frac{\exp\Bigl{(}\texttt{sim}\bigl{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{i}^{(m)})\bigr{)}/\tau\Bigr{)}}{\sum_{j=1}^{N}\sum_{{\bm{l}}_{j}^{(m)}\sim\{{\bm{l}}_{j}\}}\exp\Big{(}\texttt{sim}\bigl{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{j}^{(m^{\prime})})\bigr{)}/\tau\Big{)}},$ (2)
where $M$ is the number of regions ${\bm{l}}_{i}^{(m)}$ sampled per image. We set $M=4$ by default. We will discuss implementing the Prompter in Section 3.3, and generating ${\bm{l}}_{i}^{(m)}$ with ${\bm{y}}_{i}^{(m)}$ in Section 4. ${\mathcal{L}}_{T\rightarrow R}$ is the symmetric contrastive loss normalized along text-to-region axis, just like in Equation 1. We define ${\mathcal{L}}_{\text{{CLOC}}}=({\mathcal{L}}_{R\rightarrow T}+{\mathcal{L}}_{T\rightarrow R})/2$.
@@
${\mathcal{L}}_{\text{grounding}}:=\frac{1}{4MN}\sum_{i=1}^{N}\sum_{{\bm{l}}_{i}^{(m)}\sim\{{\bm{l}}_{i}\}}\|{\bm{l}}_{i}^{(m)}-\texttt{BoxHead}\big{(}{\bm{z}}({\bm{y}}_{i}^{(m)})\big{)}\|_{2},$ (3)
where ${\bm{z}}({\bm{y}}):=\texttt{Prompter}(f_{T}({\bm{y}}),f_{I}({\bm{x}}))$ is the grounded embedding conditioned on the text (encoded by the CLIP text encoder). The overall loss is
${\mathcal{L}}:={\mathcal{L}}_{\text{CLIP}}+\lambda({\mathcal{L}}_{\text{{CLOC}}}+{\mathcal{L}}_{\text{grounding}}),$ (4)
where $\lambda$ is a weighting scalar. In experiments, we set $\lambda$ to be the ratio of images in the mini-batch that contain region labels without extra tuning. All the learnable parameters are trained end-to-end.

### v2

@@ Abstract
CLIP has been a celebrated method for training vision encoders to generate image/text representations facilitating various applications. Recently, it has been widely adopted as the vision backbone of multimodal large language models (MLLMs). The success of CLIP relies on aligning web-crawled noisy text annotations at image levels. However, such criteria may be insufficient for downstream tasks in need of fine-grained vision representations, especially when understanding region-level is demanding for MLLMs. We improve the localization capability of CLIP with several advances. Our proposed pre-training method, Contrastive Localized Language-Image Pre-training (CLOC), complements CLIP with region-text contrastive loss and modules. We formulate a new concept, promptable embeddings, of which the encoder produces image embeddings easy to transform into region representations given spatial hints. To support large-scale pre-training, we design a visually-enriched and spatially-localized captioning framework to effectively generate region-text labels. By scaling up to billions of annotated images, CLOC enables high-quality regional embeddings for recognition and retrieval tasks, and can be a drop-in replacement of CLIP to enhance MLLMs, especially on referring and grounding tasks.
@@ 1 Introduction
Vision-language (VL) pre-training has been an important foundation for the recent tremendous growth of multimodal applications. Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021; Jia et al., 2021) is a successful VL representation learning method that connects images and text by contrastive training on large-scale data collected from the Web. Strong transferability and generalizability have been proven in extensive downstream tasks such as zero-shot image classification and image-text retrieval. Even beyond, CLIP has become arguably the default choice of vision backbone for multimodal large language models (MLLMs) (Liu et al., 2023; McKinzie et al., 2024) due to its superior prior knowledge in aligning vision and language (Tong et al., 2024).
@@
Figure 1: Our CLOC pre-training framework. (1) A visually-enriched and spatially-localized captioning pipeline pseudo-labels bounding boxes with detailed descriptions for key regions. (2) A lightweight Prompter attached on the CLIP image encoder can be prompted to transform the image embedding into region-focused features. All parameters are trained end-to-end from scratch with our contrastive localized language-image loss on the annotated region-text datasets. After pre-training, (3a) region features can be generated via the Prompter for region-text tasks like object classification in a training-free fashion. (3b) The image encoder, along with the optional Prompter, can also strengthen MLLMs fine-tuning by enhancing their fine-grained image understanding capabilities.
@@
As VL research gets increasing attention, various advanced multimodal tasks are demanding stronger vision capabilities. For instance, recent MLLMs (Rasheed et al., 2024; Ren et al., 2024; Lai et al., 2023; Chen et al., 2023; Peng et al., 2023) have been focusing on more fine-grained understanding tasks that require comprehension of the semantics at region levels such as visual question answering (VQA) with referring and grounding instructions. These MLLMs are fine-tuned on referring and grounding data with CLIP as the vision backbone, as seen in works like Kosmos-2 (Peng et al., 2023) and Ferret (You et al., 2023; Zhang et al., 2024). Due to the need for such region-level understanding, CLIP, which aligns entire images with text captions, seems insufficient, as its regular image-text contrastive loss primarily emphasizes global semantics.
@@
To remedy such core localization capability for CLIP, we ask a challenging and fundamental question: can we pre-train a stronger image encoder (1) with enhanced localization capability that can be inherently integrated into MLLMs, (2) and refine CLIP’s original image embeddings to the region level for zero-shot recognition/retrieval?
@@
Here, we explore a data-driven approach that complements the original CLIP image-text pre-training objective with explicit region-text supervision. Though conceptually simple, several challenges exist. First, it lacks public datasets with region-text annotations at scales large enough for CLIP training, which typically requires hundreds of millions even billions of images. Existing region-text corpus like Visual Genome (Krishna et al., 2017) contains about $108$K images, and the largest noisy-labeled grounded dataset GRIT (Peng et al., 2023) features only around $20$M images. Indeed, such deficiency of labeled datasets has probably limited the literature to mainly consider semi-supervised or weakly-supervised approaches as somewhat a compromise (Naeem et al., 2023; Yao et al., 2022, 2023a).
@@
  • We design a large-scale pseudo-labeling engine to support CLOC training. We combine visual-enriched image captioners and open-vocabulary detectors for an effective recipe that improves previous practice of region annotations (Minderer et al., 2024; Peng et al., 2023). This approach yields a two-billion image-text dataset with fine-grained region-text annotations, which serves as the foundation for training our CLOC model.
@@
  • We are working on releasing our pre-trained checkpoints and the constructed region-text annotations along with the final version to accelerate future research.
@@ 2 Related Work
Vision encoder pre-training.
A popular approach to MLLMs like LLaVA (Liu et al., 2023), typically connects a vision encoder (e.g., ViT (Dosovitskiy et al., 2021)) to digest visual inputs and maps them to the LLM decoder input space as token embeddings. Among various types of vision encoders (Oquab et al., 2023; He et al., 2022), CLIP (Radford et al., 2021; Jia et al., 2021) becomes the most popular choice, due to its superior performance on MLLM benchmarks reported by recent studies (Tong et al., 2024).
Other training approaches like captioning loss (Tschannen et al., 2024) are also popular, Wan et al. (2024) further incorporates bounding box coordinates. However, they need training the encoder with a decoder with smaller batch sizes thus less scalable and efficient than CLIP. Also, the vision embeddings are not directly aligned with languages thus more limited for search or retrieval tasks. Therefore, our scope focuses on improving the CLIP approach.
@@
Since CLIP was introduced, many follow-up works have been proposed to improve it from various aspects, for different target tasks, and with different approaches. From the aspect relevant to our work, improving the localization capability, most works specifically focus on the downstream dense vision tasks such as open-vocabulary detection (Minderer et al., 2024; Yao et al., 2022; Wu et al., 2023). Another less and arguably more challenging thread is to maintain the generalizability of CLIP on image-level tasks while improving localization. Recent works (Naeem et al., 2023; Bica et al., 2024; Dong et al., 2023) combine localization-enhancing unsupervised objectives with the CLIP loss, but do not attempt with supervision on large-scale explicit pseudo-labeled data like ours and are more computation overhead. Alpha-CLIP (Sun et al., 2024) shows that the SAM (Kirillov et al., 2023) can provide useful conditions for CLIP.
@@
The literature has been exploring scalable ways to generate high-quality synthetic annotations for CLIP. For instance, several works demonstrate that visually-enriched image captions improve CLIP (Lai et al., 2024). MOFI (Wu et al., 2024) augments CLIP with an extreme multi-classification task. However, these works only consider image-level annotations but not explicit region-level labels. In the context of dense vision tasks like open-vocabulary detection and segmentation, pseudo-labeling in a self-training paradigm has proven an effective approach (Kirillov et al., 2023; Minderer et al., 2024). We are inspired by these efforts and combine them to enhance CLIP’s localization capabilities. Our approach is promising, since the advance of these labeling methods can further improve ours in the future.
3 CLOC
3.1 From Image-Text to Region-Text Alignment
CLIP (Radford et al., 2021) contrastively aligns the embedding from a pair of image and text encoders ($f_{I}$ and $f_{T}$). Let a mini-batch of $N$ image-text pairs $\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ be sampled from the large-scale training set during each training iteration. The contrastive loss is defined as follows:
${\mathcal{L}}_{\text{CLIP}}$ $:=({\mathcal{L}}_{I\rightarrow T}+{\mathcal{L}}_{T\rightarrow I})/2.\>$ (1)
${\mathcal{L}}_{I\rightarrow T}$ $:=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\big{(}\texttt{sim}(f_{I}({\bm{x}}_{i}),f_{T}({\bm{y}}_{i}))/\tau\big{)}}{\sum_{j=1}^{N}\exp\big{(}\texttt{sim}(f_{I}({\bm{x}}_{i}),f_{T}({\bm{y}}_{j}))/\tau\big{)}},$
where $\texttt{sim}(\cdot,\cdot)$ is the similarity function and $\tau$ is a (learnable) temperature. The CLIP loss ${\mathcal{L}}_{\text{CLIP}}$ averages the symmetrical contrastive loss in which cross-entropy normalized along image-to-text and text-to-image axes, respectively.
Conceptually, ${\mathcal{L}}_{\text{CLIP}}$ aligns images with their associated text, but it overlooks subimage semantics. We propose augmenting this with region-text alignment on top of ${\mathcal{L}}_{\text{CLIP}}$. Specifically, assume an image-text pair $({\bm{x}},{\bm{y}})$ can be decomposed into image regions ${\bm{x}}^{(1)},\dots,{\bm{x}}^{(m)}$, and there exist regional captions ${\bm{y}}^{\prime(m)}$ that describe the corresponding regions ${\bm{x}}^{(m)}$. Thus, the original input $({\bm{x}},{\bm{y}})$ becomes region-text pairs $\{({\bm{x}}^{(1)},{\bm{y}}^{\prime(1)}),\dots,({\bm{x}}^{(m)},{\bm{y}}^{\prime(m)})\}$, and $({\bm{x}},{\bm{y}})$ is a special case when the “region” itself is the whole image. We identify several research questions and will answer them in the following sections:
@@ 3.2 Promptable Embeddings
We re-formulate the CLIP loss based on image-text pairs $({\bm{x}},{\bm{y}})$ into a localized language-image contrastive loss for region-text alignment based on triplets of ($(\{{\bm{l}}\},{\bm{x}},{\bm{y}})$, where ${\bm{l}}$ is a location representation such as a bounding box, and possibly there are several boxes as a set $\{{\bm{l}}\}$ per image. To make it compatible with CLIP training, we construct a promptable embedding transform module, or in short, region prompter ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}({\bm{x}}))$, that extracts the region embedding specified by ${\bm{l}}$ from the image embedding $f_{I}({\bm{x}})$. This formulation is inspired by the success of the segmentation model SAM (Kirillov et al., 2023) which predicts segmentation masks conditioned on location prompt (e.g., a box), while CLOC predicts a region embedding conditioned on ${\bm{l}}$ instead.
@@
${\mathcal{L}}_{R\rightarrow T}:=-\frac{1}{MN}\sum_{i=1}^{N}\sum_{{\bm{l}}_{i}^{(m)}\sim\{{\bm{l}}_{i}\}}\log{\bm{p}}$ (2)
${\bm{p}}=\frac{\exp\Bigl{(}\texttt{sim}\bigl{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{i}^{(m)})\bigr{)}/\tau\Bigr{)}}{\sum_{j=1}^{N}\sum_{{\bm{l}}_{j}^{(m)}\sim\{{\bm{l}}_{j}\}}\exp\Big{(}\texttt{sim}\bigl{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{j}^{(m^{\prime})})\bigr{)}/\tau\Big{)}},$
where $M$ is the number of regions ${\bm{l}}_{i}^{(m)}$ sampled randomly per image. We set $M=4$ by default. We will discuss implementing the Prompter in Section 3.3, and generating ${\bm{l}}_{i}^{(m)}$ with ${\bm{y}}_{i}^{(m)}$ in Section 4. ${\mathcal{L}}_{T\rightarrow R}$ is the symmetric contrastive loss normalized along text-to-region axis, just like in Equation 1. We define ${\mathcal{L}}_{\text{{CLOC}}}=({\mathcal{L}}_{R\rightarrow T}+{\mathcal{L}}_{T\rightarrow R})/2$.
@@
${\mathcal{L}}_{\text{grounding}}:=\frac{1}{4MN}\sum_{i=1}^{N}\sum_{{\bm{l}}_{i}^{(m)}\sim\{{\bm{l}}_{i}\}}\|{\bm{l}}_{i}^{(m)}-\texttt{Head}\big{(}{\bm{z}}({\bm{y}}_{i}^{(m)})\big{)}\|_{2},$ (3)
where ${\bm{z}}({\bm{y}}):=\texttt{Prompter}(f_{T}({\bm{y}}),f_{I}({\bm{x}}))$ is the grounded embedding conditioned on the text (encoded by the CLIP text encoder). All the learnable parameters are trained end-to-end. With a scalar $\lambda$, the overall loss is
${\mathcal{L}}:={\mathcal{L}}_{\text{CLIP}}+\lambda({\mathcal{L}}_{\text{{CLOC}}}+{\mathcal{L}}_{\text{grounding}}).$ (4)